# Efficient Tensor Completion Algorithms for Highly Oscillatory Operators

### v2

@@ 7 Conclusion
In this work, we proposed two low-complexity butterfly completion algorithms, respectively based on ALS and ADAM, for reconstructing highly oscillatory operators discretized as $n\times n$ matrices requiring only $\mathcal{O}(n\log n)$ observed matrix entries and $\mathcal{O}(n\log^{3}n)$ computational cost empirically. The proposed algorithms are based on the reformulation of butterfly matrices into higher order tensors, which allows for efficient implementation via tensor computations. Numerical results demonstrate the superior accuracy, convergence, and efficiency of the proposed algorithms for highly oscillatory operators such as Radon transforms and Green’s functions in seismic applications, when compared with other completion algorithms such as low-rank matrix completion and QTT completion.
Future work includes exploiting butterfly completion for stochastic low-rank decompositions, high-performance and parallel implementations of the butterfly contraction operations, as well as integrating the tensorized butterfly representation into neural network architectures to facilitate efficient learning with butterfly matrices.
@@ Appendix A Efficient Implementation of ALS for Butterfly Completion
Implementation of the ALS algorithm for butterfly completion as described in Algorithm 1 efficiently can be a challenging task when the number of levels of the butterfly decomposition increase. This is due to computations with high-order parameter tensors with their order depending on the number of levels in the butterfly decomposition. In this section, we describe a method to efficiently implement ALS for butterfly completion and QTT completion with a single software kernel that involves sparse tensor computations. This method also allows for efficient sampling of entries of a butterfly matrix (Section 3.4 of [57]).
@@ A.1 ALS for QTT completion
It should be noted that the Algorithm 5 is exactly the same as Algorithm 1 except for the indexing of the matrix-vector products used to form the left and right-hand sides. Therefore, by introducing an indexing on the input tensor entries such that the outer tensors are reshaped into matrices and the inner tensors are reshaped into order $3$ tensors, one can use exactly the same software kernel for tensor train, QTT and butterfly completion. This also means that the computational benefits of storing the sparse tensor entries in the compressed sparse fiber (CSF) format [80] to amortize the matrix-vector products involved in each solve can be translated to the butterfly completion algorithms. We describe this indexing in the next subsection.
A.2 Indexing for butterfly completion algorithms
To reshape the higher-order tensors in tensorized butterfly factorization such that the outermost tensors are represented as matrices and the inner tensors are represented as order $3$ tensors, we use the inverse of the function defined in (14). Precisely, we use a function $\psi^{-1}(\cdot)$ that takes an $p$-bit binary number as an input and outputs an integer given as
(33) $\psi^{-1}(b_{0},\dots,b_{p-1})=\sum_{i=0}^{p-1}b_{i}\,2^{i}.$
$\bm{z}^{1}\big(k_{1},:\big)$ $=\bm{s}^{1}(i_{0},\dots,i_{L},:)$
$\bm{z}^{L+2}\big(k_{L+2},:\big)$ $=\bm{s}^{L+2}(j_{0},\dots,j_{L},:)$
$\bm{Z}^{m+1}\big(k_{m+1},:,:\big)$ $=\bm{S}^{m+1}(i_{0},\dots,i_{L-m},j_{0},\dots,j_{m-1},:,:),$
Algorithm 5 ALS for QTT
@@ A.2 Indexing for butterfly completion algorithms
By using order 2 and order 3 tensor representations $\mathcal{Z}$ that adhere to the tensor train format, any tensor-train completion algorithm can be used to optimize the factors. Note that similar reshapes can be defined for the QTT format as well to utilize the implementation of tensor-train completion algorithms. We use the above reshapes for implementing Algorithm 1 and Algorithm 5 with a tensor-train completion optimization algorithm. In our implementation [^2] https://github.com/navjo2323/butterfly-matrix-completion of ALS for tensor-train completion, we transpose the sparse input tensor according to the index being solved for. This results in increasing the computational efficiency by using BLAS-$3$ instead of BLAS-$2$ operations to form the left-hand sides (see [79]) without constructing the CSF trees. The software kernel can further be optimized by using CSF trees which would allow for further increase in computational efficiency by reusing the matrix-vector products needed to form left and right-hand sides (see [81]).

### v3

@@ 7 Conclusion
In this work, we proposed three low-complexity butterfly completion algorithms, respectively based on ALS, ADAM and ADF, for reconstructing highly oscillatory operators discretized as $n\times n$ matrices requiring only $\mathcal{O}(n\log n)$ observed matrix entries and $\mathcal{O}(n\log^{3}n)$ computational cost empirically. The proposed algorithms are based on the reformulation of butterfly matrices into higher order tensors, which allows for efficient implementation via tensor computations. Numerical results demonstrate the superior accuracy, convergence, and efficiency of the proposed algorithms for highly oscillatory operators such as Radon transforms and Green’s functions in seismic applications, when compared with other completion algorithms such as low-rank matrix completion and QTT completion. While our experiments utilized only synthetic datasets, the oscillatory operators studied here, such as high-frequency Green’s functions and Radon transforms, form the mathematical foundation of several critical imaging tasks, where physical or cost constraints limit data acquisition. The proposed butterfly completion algorithm is potentially useful for addressing missing-trace problem for high-resolution reverse time migration of seismic exploration, the frequency-domain gaps inherent in low-dose medical CT, and the “missing wedge” problem in electron tomography.
Future work includes exploiting butterfly completion for stochastic low-rank decompositions, high-performance and parallel implementations of the butterfly contraction operations, as well as integrating the tensorized reformulation of the butterfly decomposition into neural network architectures to facilitate efficient learning with butterfly matrices.
@@ Appendix A Efficient Implementation of ALS for Butterfly Completion
Implementation of the ALS algorithm for butterfly completion as described in Algorithm 1 efficiently can be a challenging task when the number of levels of the butterfly decomposition increase. This is due to computations with high-order parameter tensors with their order depending on the number of levels in the butterfly decomposition. In this section, we describe a method to efficiently implement ALS for butterfly completion and QTT completion with a single software kernel that involves sparse tensor computations. This method also allows for efficient sampling of entries of a butterfly matrix (Section 3.4 of [liu2021sparse]). Our implementation is in Python and is open-source  [^2] https://github.com/navjo2323/butterfly-matrix-completion.
@@ A.1 ALS for QTT completion
It should be noted that the Algorithm 5 is exactly the same as Algorithm 1 except for the indexing of the matrix-vector products used to form the left and right-hand sides. Therefore, by introducing an indexing on the input tensor entries such that the outer tensors are reshaped into matrices and the inner tensors are reshaped into order $3$ tensors, one can use exactly the same software kernel for tensor train, QTT and butterfly completion. This also means that the computational benefits of storing the sparse tensor entries in the compressed sparse fiber (CSF) format [smith2015tensor] to amortize the matrix-vector products involved in each solve can be translated to the butterfly completion algorithms. We describe this indexing in the next subsection.
Algorithm 5 ALS for QTT Completion
@@
A.2 Indexing for butterfly completion algorithms
To reshape the higher-order tensors in tensorized butterfly factorization such that the outermost tensors are represented as matrices and the inner tensors are represented as order $3$ tensors, we use the inverse of the function defined in (14). Precisely, we use a function $\psi^{-1}(\cdot)$ that takes a $p$-bit binary number as input and outputs an integer given as
(33) $\psi^{-1}(b_{0},\dots,b_{p-1})=\sum_{i=0}^{p-1}b_{i}\,2^{p-1-i}.$
$\bm{z}^{1}\big(k_{1},:\big)$ $=\bm{s}^{1}(i_{0},\dots,i_{L},:)$
$\bm{z}^{L+2}\big(k_{L+2},:\big)$ $=\bm{s}^{L+2}(j_{0},\dots,j_{L},:)$
(34) $\bm{Z}^{m+1}\big(k_{m+1},:,:\big)$ $=\bm{S}^{m+1}(i_{0},\dots,i_{L-m},j_{0},\dots,j_{m-1},:,:),$
@@ A.2 Indexing for butterfly completion algorithms
By using order 2 and order 3 tensor representations $\mathcal{Z}$ that adhere to the tensor train format, any tensor-train completion algorithm can be used to optimize the factors. Note that similar reshapes can be defined for the QTT format as well to utilize the implementation of tensor-train completion algorithms. We use the above reshapes for implementing Algorithm 1 and Algorithm 5 with a tensor-train completion optimization algorithm. In our implementation of ALS for tensor-train completion, we transpose the sparse input tensor according to the index being solved for. This results in increasing the computational efficiency by using BLAS-$3$ instead of BLAS-$2$ operations to form the left-hand sides (see [singh2022distributed]) without constructing the CSF trees. The software kernel can further be optimized by using the computational trees which would allow for further increase in computational efficiency by reusing the matrix-vector products needed to form left and right-hand sides (see [smith2016exploration]).
A.3 Orthogonalization sweeps for butterfly ADF
[…]
QR factorization
For the outer factor $\bm{Z}^{1}$, we perform independent reduced QR factorizations of each of the $2^{L}$ blocks of size $(c\times r)$, yielding orthonormal blocks $\bm{Q}_{k_{1}}$ and upper-triangular remainders $\bm{R}_{k_{1}}$. Here recall that $k_{1}=\psi^{-1}(i_{0},\dots,i_{L})$. For an inner factor $\boldsymbol{\mathcal{Z}}^{m+1}$ with $m\in\{1,\dots,L\}$, the mode to be folded is $i_{L-m}$, since this is the outermost row index that appears in $\boldsymbol{\mathcal{Z}}^{m+1}$ but not in the next factor $\boldsymbol{\mathcal{Z}}^{m+2}$ (compare their index sets in (34)). In the encoding $k_{m+1}=\psi^{-1}(i_{0},\dots,i_{L-m},j_{0},\dots,j_{m-1})$, the index $i_{L-m}$ occupies bit position $m$ from the least significant bit. We therefore pair the $2^{L+1}$ matrices $\bm{Z}^{m+1}(k_{m+1},:,:)\in\mathbb{R}^{r\times r}$ over bit position $m$ in $k_{m+1}$: for each pair $(k^{(0)},k^{(1)})$ differing only at bit $m$ (with $k^{(0)}$ having bit $m$ equal to 0 and $k^{(1)}$ having bit $m$ equal to $1$), we concatenate along the left-rank axis to form a $(2r\times r)$ matrix and compute its reduced QR factorization. The two halves of $\bm{Q}$ are assigned back to indices $k^{(0)}$ and $k^{(1)}$, yielding $2^{L+1}$ matrices with pairwise orthonormal columns, while the remainder $\bm{R}\in\mathbb{R}^{r\times r}$ must be absorbed into the adjacent factor $\boldsymbol{\mathcal{Z}}^{m+2}$.
Absorption of $\bm{R}$
Each remainder $\bm{R}$ is left-multiplied into the matrices of $\boldsymbol{\mathcal{Z}}^{m+2}$ that share the same remaining indices. Comparing the index sets of $\boldsymbol{\mathcal{Z}}^{m+1}$ and $\boldsymbol{\mathcal{Z}}^{m+2}$, the transition from level $m+1$ to $m+2$ drops $i_{L-m}$ and introduces $j_{m}$. The new index $j_{m}$ occupies bit position 0 (the least significant bit) in $k_{m+2}=\psi^{-1}(i_{0},\dots,i_{L-m-1},j_{0},\dots,j_{m})$. Each remainder $\bm{R}$ from the QR step therefore multiplies both matrices of the pair $(k^{(0)},k^{(1)})$ differing at bit 0 in $k_{m+2}$:
(35) $\bm{Z}^{m+2}(k^{(b)},:,:)\leftarrow\bm{R}\,\bm{Z}^{m+2}(k^{(b)},:,:),\quad b\in\{0,1\}.$
When the target is the right outer factor $\bm{Z}^{L+2}$, each block $\bm{Z}^{L+2}_{k_{L+2}}$ of size $(r\times c)$ is instead updated as $\bm{Z}^{L+2}_{k_{L+2}}\leftarrow\bm{R}_{k_{L+2}}\,\bm{Z}^{L+2}_{k_{L+2}}$.
The complete ADF algorithm for butterfly matrix completion is summarized in Algorithm 6. The gradient computation is the same as that used in Algorithm 3 and the optimal step size for each slice of the computed tensors are computed as suggested in [grasedyck2015variants].
Algorithm 6 ADF for Butterfly Completion
1:Input: Observed matrix $\bm{T}_{\Omega}$ (or its tensorization $\boldsymbol{\mathcal{T}}_{\Omega}$), rank $r$, initial guess $\boldsymbol{\mathcal{S}}^{i},i=1,\ldots,L+2$, max iterations $t_{\max}$, and convergence threshold $\epsilon$.
[…]
5:for $\text{iter}=1,\dots,t_{\max}$ do
6:  for $l=1,2,\dots,L+2$ do
7:   $\boldsymbol{\mathcal{N}}\leftarrow\textsc{Gradient}(\boldsymbol{\mathcal{G}},\Omega,\boldsymbol{\mathcal{Z}},l,L)$ $\triangleright$ As in Algorithm 3
8:   $\boldsymbol{\mathcal{Z}}\leftarrow\textsc{Reconstruct}_{\Omega}(\boldsymbol{\mathcal{Z}}|_{\boldsymbol{\mathcal{Z}}^{l}\leftarrow\boldsymbol{\mathcal{N}}})$
9:   $\alpha_{s}\leftarrow\|\boldsymbol{\mathcal{N}}_{(s)}\|_{F}^{2}\,/\,\|\bm{z}\|_{2}^{2}$ $\triangleright$ Overrelaxation for each slice as in [grasedyck2015variants]
10:   $\boldsymbol{\mathcal{Z}}^{l}_{s}\leftarrow\boldsymbol{\mathcal{Z}}^{l}_{s}+\alpha_{s}\,\boldsymbol{\mathcal{N}}_{s}$ $\triangleright$ Slice-wise update
11:   $\boldsymbol{\mathcal{G}}_{s}\leftarrow\boldsymbol{\mathcal{G}}_{s}-\alpha_{s}\,\boldsymbol{\mathcal{Z}}_{s}$ $\triangleright$ Slice-wise update
12:   if $l\leq L+1$ then $\triangleright$ Batched QR and absorb into next level
13:     $\boldsymbol{\mathcal{Z}}^{l},\bm{R}\leftarrow\textsc{QR}(\boldsymbol{\mathcal{Z}}^{l},L,l)$
14:     $\boldsymbol{\mathcal{Z}}^{l+1}\leftarrow\textsc{Absorb}(\bm{R},\boldsymbol{\mathcal{Z}}^{l+1},L,l)$
15:   end if
16:  end for
17:  if $\|\bm{g}\|_{2}/\|\boldsymbol{\mathcal{T}}_{\Omega}\|_{F}<\epsilon$ then
18:   break
19:  end if
20:end for